zarkam21:

True or False: For a trigonometric function, y = f(x), then x = F-1(y). Explain your answer.

SmokeyBrown:

If I'm interpreting this correctly, the statement is saying that in a trigonometric function if y is a certain value for a certain value of x, then x will always be a certain value for a certain value of y.

SmokeyBrown:

Now, thinking back to trig functions you've seen before, can you find any examples to support or disprove this statement?

zarkam21:

I guess, I am not really understanding this

SmokeyBrown:

Ok, so there's a test you can do to see if this statement is true or not. If you draw a horizontal line through the function, if the statement is true, the horizontal line should only pass through the curve of the function once. If the horizontal line passes through the function more than once, you know that any particular y value can match with many different x values, which disproves the statement.

zarkam21:

It is passed once , it think :/

zarkam21:

SO true because it passed the horizontal line test and pass through the curve of the function once

SmokeyBrown:

Ok, just checking, what function did you use the horizontal line test on?

zarkam21:

first one

zarkam21:

y = f(x),

SmokeyBrown:

Ah, gotcha. So, y=f(x) isn't really a specific function. It's just defining what a function needs to be: for every value of x in the domain there is 1 and only 1 value of y. You're going to want to test a specific function, like, particularly a trigonometric equation. These are equations that use trigonometric terms, like sine, cosine and tangent. You could use the cosine function from the question we did earlier, for instance.

zarkam21:

Okay so y = cosx

SmokeyBrown:

Sure. How does the horizontal line test work for y = cos(x)?

zarkam21:

nope it fails it

SmokeyBrown:

Great! So that means we've disproved the statement.

zarkam21:

So false, because it does not pass the horizontal line test. It passes through the curve of the function more than once

SmokeyBrown:

Yup. Sounds pretty good

zarkam21:

are you going to bed now?

SmokeyBrown:

Nah, I'll be up for a while.

zarkam21:

True or False: For a one-to-one function, y = f(x), then x = f-1(y). Explain your answer.

SmokeyBrown:

Ok, this one is pretty easy to explain. A one-to-one function, by definition, has one y value for every y value and one x value for every y value. No more, no less.

zarkam21:

So true

zarkam21:

and the explanation part is the defnition of what a one to one function is

SmokeyBrown:

Yup, you got it

SmokeyBrown:

I think that'd work, yes

zarkam21:

True or False: For any function, x = f-1(y), then y = f(x). Explain your answer.

SmokeyBrown:

For this one, you can think of it as similar to the first question in this series, only instead of "every y-value only has one x value" it's "every x value has only one y-value:"

zarkam21:

So it owld be the same answer right

SmokeyBrown:

By the definition of a function this has to be true. You can do the vertical line test to demonstrate this, but you know enough about functions to accept it at face value, I think.

zarkam21:

Except we would use the vertical line test

SmokeyBrown:

Yup

SmokeyBrown:

Oh, and you can use any function, not just trig functions

zarkam21:

True, because it passes the vertical line test. It does not pass through the curve of the function more than once.

SmokeyBrown:

I completely agree.

zarkam21:

1 attachment
SmokeyBrown:

We can actually take a look at a the sine function to see whether this one is true or not

1 attachment
zarkam21:

no it is not

zarkam21:

because it fails the horizontal line test

zarkam21:

or wait its the vertical line test right

SmokeyBrown:

So, this one is a little different from the others. It's asking about the direction of the line over a certain interval. So, between the two intervals mentioned, pi/2 and 3pi/2, does the function increase or decrease?

zarkam21:

increase

SmokeyBrown:

Yes, you got it!

zarkam21:

So true, because the direction of the lines over the two certain intervals are increasing within each other

SmokeyBrown:

Exactly

zarkam21:

Going to bed now??

SmokeyBrown:

Heheh, no, I don't have classes until late tomorrow. College is weird like that.

zarkam21:

Lol i bet

zarkam21:

Explain the meaning of y = cos^-1x.

SmokeyBrown:

So, whenever anything is taken to the power of negative 1, it means you are getting the reciprocal of that value. For instance, 2^(-1) is the same as 1/2. 5^(-1) is the same as 1/5. in the same way, Cos(x)^-1 is the same as 1/cos(x)

SmokeyBrown:

It's difficult to tell with the notation here, but I think that is what you meant?

zarkam21:

1 attachment
zarkam21:

I screenshot it so you can see

SmokeyBrown:

Ok, so I did understand correctly. That's good. I stand by what I said.

zarkam21:

Okay so the meaning of y = cos^-1x is whenever anything is taken to the power of negative 1, it means you are getting the reciprocal of that value. From this function given, we can see that Cos(x)^-1 is the same as 1/cos(x)

SmokeyBrown:

Yup, I think you understand.

zarkam21:

1 attachment
zarkam21:

Thsi would be true

zarkam21:

If i remember correctly, when a function is undefined, it is undefined on an interval or at certain points. whether and/or where a function is undefined depends on the given domain of the function.

SmokeyBrown:

That's interesting. Keep in mind that sec(x) is equal to 1/cos(x). So, Sec(x)^-1, would be 1/sec(x), which is actually the same as cos(x). I would say that cos(0.5) is defined, so I would say the statement is false.

zarkam21:

could you check this over

zarkam21:

1 attachment
SmokeyBrown:

I'd say your math checks out. And I also want to make a correction to what I said earlier. With plain numbers, what I said about something to the power of (-1) is true. In this case, it means you're taking the inverse of the function. You can think of this as moving backwards; usually, you'd convert x, the angle or arc length, to y, the value of, say, cosine for that particular value. With an inverse function, you're given the value of cosine and use the function to find the value of the angle or, in this case the arc length. Even with my wrong explanation, I'm impressed you got these right!

SmokeyBrown:

I think now is when I'm going to start going to sleep. Anything else before I head out?

zarkam21:

Nope, thats all Just to make sure even with your wrong explanation is there any specific question I need to correct, or are my answers still correct

SmokeyBrown:

Nah, you got it right

zarkam21:

Nvm, you can go to bed we can continue tomorrow or something. I don't want to hold you up

zarkam21:

Goodnight, and thank you for all your help.. I really do appreciate it

SmokeyBrown:

No worries. See you around!